OpenStudy (anonymous):

Which of the following equations could be used to solve the given equation? 9x + 26 + 7x - 17 = 2x + (-3x) + 5x

OpenStudy (anonymous):

What are you asking here?? :P ^.^

OpenStudy (anonymous):

Welcome to OS! We need to see the "following equations" if you can post them, please.

OpenStudy (anonymous):

sorry it didnt paste the choices. a)16x+9=0 b)16x+9=4x c)16x+11=4x

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Okay first we need to simplify your original equation: 9x + 26 + 7x - 17 = 2x + (-3x) + 5x 16x + 9 = 4x That looks familiar.

OpenStudy (anonymous):

Basically just simplify your terms by combining like terms (like 2x & 4x are like terms, because both are [some constant] x [some variable]

OpenStudy (anonymous):

And the variable is the same, x

OpenStudy (anonymous):

Thankyou both so much.

OpenStudy (dan815):

D